OpenStudy (aravindg):

If f(x)= x2 + 3 x + 2, find that value of x for which f(x + 1) = f(x).

OpenStudy (anonymous):

find the vertex then minus a half

OpenStudy (anonymous):

x=-b/(2a) x=-3/2 so the points u want are x=-3/2-1/2=-4/2=-2 and x=-3/2+1/2=-2/2=-1

OpenStudy (anonymous):

x^2+3x+2=(x+2)(x+1) so when x=-1 or -2, y=0 so it is right

OpenStudy (aravindg):

thx man

OpenStudy (anonymous):

no prob